15-th Conference of the International Linear Algebra Society

April 1, 2008

Professor Assaf Goldberger

Tel Aviv University Tel Aviv, Israel

Dear Professor Goldberger:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society invites you to participate in our Conference delivering a lecture about your recent research results in the Contributed Papers Session.

The 15-th Conference will be held in Cancún, México, from June 15 to June 21, 2008. Some useful information is available at the conference web-page

http://star.izt.uam.mx/ILAS08

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially

Dr. Luis Verde-Star Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa Apartado 55-534, Mexico DF 09340, Mexico Tel. +52 55 5804 4961, Fax +52 55 5804 4653 e-mail: verde@star.izt.uam.mx

April 1, 2008

Professor Wei Cheng

National University of Defense Technology Changsha, P.R.China

Dear Professor Cheng:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society invites you to participate in our Conference delivering a lecture about your recent research results in the Contributed Papers Session.

The 15-th Conference will be held in Cancún, México, from June 15 to June 21, 2008. Some useful information is available at the conference web-page

http://star.izt.uam.mx/ILAS08

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially

Dr. Luis Verde-Star Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa Apartado 55-534, Mexico DF 09340, Mexico Tel. +52 55 5804 4961, Fax +52 55 5804 4653 e-mail: verde@star.izt.uam.mx

April 1, 2008

Professor Ljiljana Cvetkovic

Faculty of Science Novi Sad, Serbia

Dear Professor Cvetkovic:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society invites you to participate in our Conference delivering a lecture about your recent research results in the Contributed Papers Session.

The 15-th Conference will be held in Cancún, México, from June 15 to June 21, 2008. Some useful information is available at the conference web-page

http://star.izt.uam.mx/ILAS08

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially

Dr. Luis Verde-Star Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa Apartado 55-534, Mexico DF 09340, Mexico Tel. +52 55 5804 4961, Fax +52 55 5804 4653 e-mail: verde@star.izt.uam.mx

April 1, 2008

Professor Vladimir Kostic

Faculty of Science Novi Sad, Serbia

Dear Professor Kostic:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society invites you to participate in our Conference delivering a lecture about your recent research results in the Contributed Papers Session.

The 15-th Conference will be held in Cancún, México, from June 15 to June 21, 2008. Some useful information is available at the conference web-page

http://star.izt.uam.mx/ILAS08

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially

Dr. Luis Verde-Star Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa Apartado 55-534, Mexico DF 09340, Mexico Tel. +52 55 5804 4961, Fax +52 55 5804 4653 e-mail: verde@star.izt.uam.mx

April 1, 2008

Professor Mohammad Shahryari

Tabriz University Tabriz, Iran

Dear Professor Shahryari:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society invites you to participate in our Conference delivering a lecture about your recent research results in the Contributed Papers Session.

The 15-th Conference will be held in Cancún, México, from June 15 to June 21, 2008. Some useful information is available at the conference web-page

http://star.izt.uam.mx/ILAS08

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially

Dr. Luis Verde-Star Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa Apartado 55-534, Mexico DF 09340, Mexico Tel. +52 55 5804 4961, Fax +52 55 5804 4653 e-mail: verde@star.izt.uam.mx

April 1, 2008

Professor Biswajit Deb

Sikkim Manipal Institute of Technology Majitar, Sikkim India

Dear Professor Deb:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society invites you to participate in our Conference delivering a lecture about your recent research results in the Contributed Papers Session.

The 15-th Conference will be held in Cancún, México, from June 15 to June 21, 2008. Some useful information is available at the conference web-page

http://star.izt.uam.mx/ILAS08

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially

Dr. Luis Verde-Star Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa Apartado 55-534, Mexico DF 09340, Mexico Tel. +52 55 5804 4961, Fax +52 55 5804 4653 e-mail: verde@star.izt.uam.mx

April 1, 2008

Professor Hossein Moosaei

Guilan, Rasht

Dear Professor Moosaei:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society invites you to participate in our Conference delivering a lecture about your recent research results in the Iran

The 15-th Conference will be held in Cancún, México, from June 15 to June 21, 2008. Some useful information is available at the conference web-page

http://star.izt.uam.mx/ILAS08

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially

Dr. Luis Verde-Star Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa Apartado 55-534, Mexico DF 09340, Mexico Tel. +52 55 5804 4961, Fax +52 55 5804 4653 e-mail: verde@star.izt.uam.mx

April 1, 2008

Professor Fanja Rakotondrajao

University of Antananarivo Antananarivo, Madagascar

Dear Professor Rakotondrajao:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society invites you to participate in our Conference delivering a lecture about your recent research results in the Contributed Papers Session.

The 15-th Conference will be held in Cancún, México, from June 15 to June 21, 2008. Some useful information is available at the conference web-page

http://star.izt.uam.mx/ILAS08

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially

Dr. Luis Verde-Star Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa Apartado 55-534, Mexico DF 09340, Mexico Tel. +52 55 5804 4961, Fax +52 55 5804 4653 e-mail: verde@star.izt.uam.mx

April 1, 2008

Professor Marie Sangare

Institut de polytechnique Fianarantsoa, Madagascar

Dear Professor Sangare:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society invites you to participate in our Conference delivering a lecture about your recent research results in the Contributed Papers Session.

The 15-th Conference will be held in Cancún, México, from June 15 to June 21, 2008. Some useful information is available at the conference web-page

http://star.izt.uam.mx/ILAS08

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially

Dr. Luis Verde-Star Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa Apartado 55-534, Mexico DF 09340, Mexico Tel. +52 55 5804 4961, Fax +52 55 5804 4653 e-mail: verde@star.izt.uam.mx

April 1, 2008

Professor Pei Yuan Wu

National Chiao Tung University Hsinchu, Taiwan

Dear Professor Wu:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society invites you to participate in our Conference delivering a lecture about your recent research results in the Contributed Papers Session.

The 15-th Conference will be held in Cancún, México, from June 15 to June 21, 2008. Some useful information is available at the conference web-page

http://star.izt.uam.mx/ILAS08

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially

Dr. Luis Verde-Star Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa Apartado 55-534, Mexico DF 09340, Mexico Tel. +52 55 5804 4961, Fax +52 55 5804 4653 e-mail: verde@star.izt.uam.mx